OpenStudy (anonymous):

derivative of ln(x^2) I know we have to use chain rule but how

OpenStudy (anonymous):

@Hero

OpenStudy (imstuck):

\[\frac{ d }{ dx }\ln(x)^{2}\]Is that it?

OpenStudy (anonymous):

2/ x ?

OpenStudy (imstuck):

The derivative of ln(x^2) is 1/ln(x^2) times the derivative of x^2, which is 2x. I get that equals\[\frac{ 2x }{ \ln(x ^{2}) }\]

OpenStudy (imstuck):

If the x is squared inside the parenthesis, then it changes the derivative.

OpenStudy (anonymous):

inside the paranthesis

OpenStudy (anonymous):

i am getting 2/x

OpenStudy (anonymous):

is it incorrect then

OpenStudy (imstuck):

Show me what you are doing.

OpenStudy (anonymous):